Circular No. 6532

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1997 A1 E. F. Helin, Jet Propulsion Laboratory, on behalf of the Near-Earth-

Asteroid Tracking Team (cf. IAUC 6341), reports the discovery of a comet on images obtained by the on-site USAF/Ground-based Electro-Optical Deep Space Surveillance Team in Hawaii. Available precise positions:

1997 UT  α_{2000}  δ_{2000}  m_{1} Observer
Jan. 10.51495 9^{h}39^{m}18.^{s}96 +16^{o}26′18′′.9 18.6 Helin
10.52334 9 39 17.97 +16 26 28.9 18.8  _{”}
10.53174 9 39 17.04 +16 26 38.6 18.9  _{”}
11.33935 9 37 46.18 +16 42 31.0 18.9 Rabinowitz
11.34959 9 37 45.11 +16 42 43.9  _{”}

E. F. Helin, K. J. Lawrence and S. Pravdo (Haleakala-NEAT/GEODSS).

1-m f/2.2 Ritchey-Chrétien + CCD. Images clearly diffuse. D. L. Rabinowitz (Table Mountain Observatory). 0.6-m reflector + CCD.

Near limiting magnitude of exposure. 15′′ coma, extending asymmetrically to the south.

NOVA CRUCIS 1996 M. Della Valle, Universitá di Padova; and S. Benetti, European South-

ern Observatory, report: “Analysis of a fully-reduced spectrum (range 375– 690 nm, resolution ∼ 1.6 nm) obtained on Jan. 10.3 UT with the ESO 3.6-m telescope + EFOSC1 confirms this object (cf. IAUC 6463, 6466, 6468) to be a galactic nova caught during the nebular phase. The spectrum is dominated by the Balmer emission lines, [O III] 500.7 and 495.9 nm, He II 468.6 nm and [Ne III] 386.9 and 396.8 nm. In particular, we note that the He II emission line is almost four times stronger than Hβ, suggesting that this nova belongs to the He/N class (Williams 1992, A.J. 104, 725). The expansion velocity provided by the FWHM of Hα and Hβ is ∼ 2000 km/s.”

SUPERNOVA 1997A IN ANONYMOUS GALAXY Benetti and Della Valle also report: “Analysis of a spectrum obtained

as above on Jan. 10.2 UT shows that supernova 1997A (cf. IAUC 6531) resembles the type Ia supernova 1994D (cf. IAUC 5946, etc.) ∼ 7–10 days past maximum. The spectrum exhibits strong Fe II and fading S II lines.”

1997 January 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden